# Minquan Cheng, Huimei Wei, Kai Wan, and Giuseppe Caire

### v2

@@ II PRELIMINARIES
In this section, we will introduce a coded caching system, placement delivery array and their relationship respectively.
@@ II-A Coded caching system
A $(K,M,N)$ coded caching system illustrated in Figure 1 contains a server storing $N$ equal-sized files $\mathcal{W}=\{W_{n}\ |\ n\in[N]\}$, $K$ users each of which can cache at most $M$ files where $0\leq M\leq N$. The sever connects the users over an error-free shared-link. An $F$-division $(K,M,N)$ coded caching scheme operates in two phases:
@@
  • Placement phase: The server divides each file into $F$ equal sized packets, i.e., $W_{n}=\{W_{n,j}\ |\ j\in[F]\}$ where $n\in[N]$, and places some packets into each user’s cache without knowing any information about future demands. Let $\mathcal{Z}_{k}$ be the packets cached by user $k$. This placement strategy is called uncoded placement, otherwise it is called coded placement.
  • Delivery phase: Each user requests one file from $\mathcal{W}$ randomly. Denote the requested file number by ${\bf d}$=$(d_{1},d_{2},\ldots,d_{K})$, i.e., user $k$ requests file $W_{d_{k}}$, where $k\in\mathcal{K},d_{k}\in[N]$. After receiving the request vector ${\bf d}$, the sever broadcasts XOR of packets with a size of at most $R_{\bf d}F$ to the users such that each user is able to decode its requested file.
@@
$R=\max\{R_{\bf d}\ |\ {\bf d}\in[N]^{K}\}$ (1)
The first well-known scheme was proposed in [2], which is referred to as the MN scheme, has the minimum load under uncoded placement [3, 4, 5, 6] but has large subpacketization which increases exponentially with the number of users.
@@ II-B Placement delivery array
In order to study the scheme with lower subpacketization, the authors in [9] proposed an interesting combinatorial structure called placement delivery array defined as follows.
@@
A PDA is called $g$-regular if each integer occurs exactly $g$ times. One can check that the following array $\mathbf{P}$ is a $2$-$(4,4,2,4)$ PDA.
@@
There are many constructions based on PDA [8, 10, 11, 12, 38, 16, 9, 24, 22, 39]. The authors in [9] showed that the MN scheme can also be represented by a PDA. Some other constructions could be also represented by appropriate PDAs such as the caching schemes based on liner block code [21], projective space [26], and combinatorial designs [27], etc. We list their performances in Table I. In addition, given a PDA, we can obtain a PDA with any number of users by the following result.
@@ III Non-half-sum disjoint packing
Recall that Condition C3333-a) of Definition 1 is called the Latin property in combinatorial design theory. Naturally, we can construct a PDA with $K=F$ by first constructing a Latin square and then putting stars in some integer entries such that the resulting array satisfies Conditions C1111, C2222, and C3333. Then we obtain the desired PDA with linear subpacketization. Specifically, we use the classical and simple construction of the Latin square $\mathbf{L}=(l_{f,k})_{f,k\in\mathbb{Z}_{v}}$ where the entry $l_{f,k}=f+k$ for each $f$, $k\in\mathbb{Z}_{v}$. Recall that $\mathbb{Z}_{v}$ is a positive integer ring, i.e., the arithmetic operations in the ring $\mathbb{Z}_{v}$. When we cyclically replace the integer entries by stars, Condition C3333 of Definition 1 is transformed into the following novel combinatorial structure for the integer entries of the first row of the Latin square.
@@
  • For each $\mathcal{D}\in\mathfrak{D}$, the half-sum of any two different elements in $\mathcal{D}$ (i.e., the sum of the two elements divided by $2$) does not appear in any block of $\mathfrak{D}$ [^1] In an NHSDP $(\mathbb{Z}_{v},\mathfrak{D})$, each element of $\mathfrak{D}$ is called block.
@@
It is not difficult to check that each column of $\mathbf{P}$ has exactly $Z=v-bg=15-2\times 4=7$ stars, there are exactly $S=30$ vectors occur in $\mathbf{P}$, and each vector occurs at most once in each row and each column. So the conditions C1111, C2222, and C3333-a) of Definition 1 hold. Finally, let us consider Condition C3333-b). Let us first consider the entries $p_{1,0}=p_{0,1}=(1,1)$. We have $p_{1,1}=p_{0,0}=*$ which satisfies Condition C3333-b). The reason for this is as follows. When $f=1$, if $k=0$ we have $k-f=-1=14\in\mathcal{D}_{1}$ in (12) and $1+0=1$, then we set $p_{1,0}=(1,1)$; if $k=1$, we have $k-f=0$, which does not appear in either $\mathcal{D}_{1}$ or $\mathcal{D}_{2}$. Thus we set $p_{1,1}=*$. Similarly we can check that all the vectors in $\mathbf{P}$ satisfy Condition C3333-b) of Definition 1. So $\mathbf{P}$ is a $(15,15,7,30)$ PDA which can realize a coded caching scheme with the memory ratio $\frac{M}{N}=\frac{7}{15}$, the subpacketization $F=K=15$, and load $R=2$. $\square$
In Example 2, each column of $\mathbf{P}$ has exactly $|\mathcal{D}_{1}|+|\mathcal{D}_{2}|=|\mathcal{D}_{1}\cup\mathcal{D}_{2}|=8$. This implies that Condition C1111 of Definition 1 is guaranteed by the first condition of NHSDP. From (13), the half-sum set is =$\mathcal{H}=\{0,3,6,7,8,9,12\}$. We can see that in Example 2 the stars of $\mathbf{P}$ in (15) are exactly the entries $p_{i,i+h}$ for each $i\in\mathbb{Z}_{15}$ and $h\in\mathcal{H}$. So the half-sums of all the blocks generates the star positions. This implies that Condition C3333-b) of Definition 1 is ensured by the second condition of NHSDP. Recall that our construction is based on the constructing a Latin square. So the condition C3333-a) always holds. Then, for any parameters $v$, $g$ and $b$, if there exists a $(v,g,b)$ NHSDP by Construction 1, we can also obtain a PDA in the following theorem, whose proof is given in Appendix A.
@@ IV Constructions of PDAs via NHSDPs
For instance, the integer $31$ in $\mathcal{D}_{(1,1,1)}$ can be obtained by $1+5+25=31$. It is evident that the number of all the possibilities of $\alpha_{1},\alpha_{2}$, and $\alpha_{3}$ is $82^{3}=$ 8, so the block $\mathcal{D}_{(1,1,1)}$ contains $8$ different integers. Similarly, we can obtain the following blocks.
$\begin{split}\mathfrak{D}=\{&\mathcal{D}_{(1,1,1)}=\{31,21,29,19,-19,-29,-21,-31\},\ \ \mathcal{D}_{(2,1,1)}=\{32,22,28,18,-18,-28,-22,-32\},\\ &\mathcal{D}_{(1,2,1)}=\{36,16,34,14,-14,-34,-16,-36\},\ \ \mathcal{D}_{(2,2,1)}=\{37,17,33,13,-13,-33,-17,-37\},\\ &\mathcal{D}_{(1,1,2)}=\{56,46,54,44,-44,-54,-46,-56\},\ \ \mathcal{D}_{(2,1,2)}=\{57,47,53,43,-43,-53,-47,-57\},\\ &\mathcal{D}_{(1,2,2)}=\{61,41,59,39,-39,-59,-41,-61\},\ \ \mathcal{D}_{(2,2,2)}=\{62,42,58,38,-38,-58,-42,-62\}\}.\end{split}$ (20)
@@
By the above equation, $\mathfrak{D}$ has $m_{1}m_{2}m_{3}=8$ blocks and the intersection of any blocks is empty. The minimum and maximum values of $\mathfrak{D}$ are $-62$ and $62$ respectively. To ensure that these two sets $[-62:0]$ and $[1:62]$ have not overlap in $\mathbb{Z}_{v}$, the value of $v$ must satisfy $v\geq 2\times 62+1=125$, since 00 is also in $\mathbb{Z}_{v}$. Next, let we verify that $(\mathbb{Z}_{125},\mathfrak{D})$ is an $(125,8,8)$ NHSDP. Now let we consider the half-sum of $31$ and $21$, i.e., $\frac{31+21}{2}=26$ which is not in any block of $\mathfrak{D}$. Similarly we can all the half-sum of any two different integers in each block of $\mathfrak{D}$.
The intuition that $(\mathbb{Z}_{125},\mathfrak{D})$ is an NHSDP lies in the fact that each the integers in $\mathfrak{D}$ can be uniquely represented by a combination $\mathcal{X}$ with some coefficient from $\mathcal{C}=\{\pm 1,\pm 2\}^{3}$ and each half-sum can only be represented by $\mathcal{X}$ with some coefficient from $\mathcal{C}^{\prime}=[-2:2]^{3}\setminus\mathcal{C}$. Clearly $\mathcal{C}\cap\mathcal{C}^{\prime}=\emptyset$ which implies that each half-sum does not occurs in $\mathcal{D}$. For instance, the integers $31$ and $21$ from $\mathcal{D}_{1}$ can be uniquely represented as
$31=1\cdot 1+1\cdot 5+1\cdot 25\ \ \text{and}\ \ \ 21=1\cdot 1+(-1)\cdot 5+1\cdot 25$
Their half-sum $26$ can be represented by
$\frac{(1\cdot 1+1\cdot 5+1\cdot 25)+(1\cdot 1+(-1)\cdot 5+1\cdot 25)}{2}=\frac{1+1}{2}\cdot 1+\frac{1-1}{2}\cdot 5+\frac{1+1}{2}\cdot 25=1\cdot 1+0\cdot 5+1\cdot 25=26.$
@@
  • The WCWL scheme: When $K=125$ and $t=61$ in [14] we have a WCWL scheme with $K=125$, the subpacketization $F=125$, and the transmission load $\frac{K-t}{2\lfloor\frac{K}{K-t+1}\rfloor}=\frac{64}{3}\approx 21.333>8$. Clearly our scheme has a lower transmission load while maintaining a lower memory ratio and the same subpacketization;
  • The CWZW scheme: In the case of $q=2$ and $m=7$, the scheme in [12] includes the schemes in [9, 22, 21]. So we only need consider the CWZW scheme. First we can obtain a CWZW scheme with $K=16$, the memory ratio $M/N=0.5>\frac{61}{125}$, the subpacketization $F^{\prime}=2^{7}=128>125$, and the transmission load $1$. Using the grouping method in Lemma 2, we have a grouping CWZW scheme with $K=125$, $M/N=0.5>\frac{61}{125}$, the subpacketization $F_{1}=\frac{16}{\text{gcd}(125,16)}\cdot F^{\prime}=2048>125$, and the transmission load $\frac{125}{16}=7.8125$. We can see that our scheme has a lower memory ratio and a smaller subpacketization than the scheme in [12], while albeit with a slight increase in the transmission load;
  • The MN scheme: An exhaustive computer search shows that we can obtain a $(125,M,N)$ grouping MN coded caching scheme with $M/N=0.5$, an appropriate subpacketization (${10\choose 5}=252$, and the transmission load $R=\frac{5}{6}\times\frac{125}{10}=\frac{125}{12}>10>8$ by using a $(10,M,N)$ MN scheme based on the grouping method in Lemma 2. It is evident that all the memory ratio, the subpacketization, and the transmission load are larger than our proposed scheme.
@@
From Construction 2, for any $n$ and $m_{1}$, $m_{2}$, $\ldots$, $m_{n}$ are positive integers, suppose the vector $\mathbf{a}=(a_{1},a_{2},\ldots,a_{n})=(m_{1},m_{2},\ldots,m_{n})\in\mathcal{A}$. The corresponding block $\mathcal{D}_{\mathbf{a}}$ is given by
@@
which contains the minimum value $-f(1)-f(2)-\dots-f(n)$ and the maximum value $f(1)+f(2)+\ldots+f(n)$. To ensure that $[-f(1)-f(2)-\dots-f(n):0]$ and $[1:f(1)+f(2)+\ldots+f(n)]$ can not overlap in $\mathbb{Z}_{v}$, the value of $v$ must satisfy
@@
Next given the parameters $v$ and $n$ (recall that $K=v$ and $g=2^{n}$), we consider the selection of $m_{1},\ldots,m_{n}$, such that the memory ratio is minimum. In other words, fixing the coded caching gain, we aim to search the coded caching scheme requiring the minimum memory ratio. Considering the constraint $v\geq 2\phi(m_{1},m_{2},\ldots,m_{n})+1$ into the optimization problem for the selection of $m_{1},m_{2},\ldots,m_{n}$, we have
@@
Problem 1 is an integer programming problem, which is NP-hard. We first relax the constraint $m_{1},m_{2},\ldots,m_{n}\in\mathbb{Z}^{+}$ to $m_{1},m_{2},\ldots,m_{n}\in\mathbb{R}^{+}$, to simplify the problem to a convex optimization problem, which is then solved by the Lagrange Multiplier Method. Finally we take the floor operation to the solution. For the ease of notation, we define that $q:=K^{1/n}$. The following theorem introduces the sub-optimal solution by the above optimization strategy, whose proof is given in Appendix C.
@@
which can realize a coded caching scheme with the memory ratio $\frac{M}{N}=1-\left(\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}}{q^{n}}\right)$ and the transmission load $R={\left\lfloor\frac{q-1}{2}\right\rfloor}^{n}$.
@@
By Theorem 3, there are $\left\lfloor\log_{3}K\right\rfloor$ memory-rate ratio points for each user number $K$. In addition, when $q$ equals to a positive odd integer such that $q\geq 3$, the solution in (23) is an optimal solution to Problem 1. Then by Theorem 3 we have the following result.

### v3

@@ II PRELIMINARIES
In this section, we will review a coded caching system, placement delivery array, and their relationship.
@@ II-A Coded caching system
A $(K,M,N)$ coded caching system illustrated in Figure 1 contains a server storing $N$ equal-sized files $\mathcal{W}=\{W_{n}\ |\ n\in[N]\}$, $K$ users each of which can cache at most $M$ files where $0\leq M\leq N$. The server connects to the users over an error-free shared-link. An $F$-division $(K,M,N)$ coded caching scheme operates in two phases:
@@
  • Placement phase: The server divides each file into $F$ equal-sized packets, i.e., $W_{n}=\{W_{n,j}\ |\ j\in[F]\}$ where $n\in[N]$, and places some packets into each user’s cache without knowing any information about future demands. Let $\mathcal{Z}_{k}$ be the packets cached by user $k$. This placement strategy is called uncoded placement, otherwise it is called coded placement.
  • Delivery phase: Each user requests one file from $\mathcal{W}$ randomly. Denote the requested file number by ${\bf d}$=$(d_{1},d_{2},\ldots,d_{K})$, i.e., user $k$ requests file $W_{d_{k}}$, where $k\in\mathcal{K},d_{k}\in[N]$. After receiving the request vector ${\bf d}$, the sever broadcasts XOR of coded packets with a size of at most $R_{\bf d}F$ to the users such that each user is able to decode its requested file.
@@
$R=\max\{R_{\bf d}\ |\ {\bf d}\in[N]^{K}\}.$ (1)
The first well-known scheme was proposed in [2], which is referred to as the MN scheme, has the minimum load under uncoded placement [3, 4, 5, 6], but the MN scheme has large subpacketization which increases exponentially with the number of users.
@@ II-B Placement delivery array
In order to study linear schemes with lower subpacketization, the authors in [9] proposed an interesting combinatorial structure called placement delivery array defined as follows.
@@
A PDA is called $g$-regular if each integer occurs exactly $g$ times. For example, one can check that the following array $\mathbf{P}$ is a $2$-$(4,4,2,4)$ PDA.
@@
In plain words, the relationship between a PDA and the corresponding coded caching scheme is as follows. Each $k$-th column of the PDA describes the cache configuration of the corresponding $k$-th user. Each file is split into $F$ packets. User $k$ caches the packets corresponding to the * symbol in the $k$-th PDA column. Thus, the amount of data cached by user $k$ is $Z$ (the number of * symbols) times the (normalized) size of one packet, which is $1/F$. This yields a memory ratio $M/N=Z/F$. Then, each integer $s\in[S]$ corresponds to the transmission of one linearly coded packet. At transmission step $s$, the severs takes all the packets of the demanded files marked by entries equal to $s$ in $\mathbf{P}$, and transmits the XOR. By condition C2 above, each XOR is formed by packets needed by (at least one) user, and present in the caches of all other users whose requests participate in the XOR. Therefore, each user obtains either $1$ or 00 packet from each XOR. It takes at most $S$ rounds to satisfy all users’ demands. In particular, there is at least one configuration of the user demands that requires $S$ rounds. The amount of transmitted data (normalized to the size of one file) is therefore $R=S/F$.
@@
There are many constructions based on PDA [8, 10, 11, 12, 38, 16, 9, 24, 22, 39]. The authors in [9] have shown that the MN scheme can also be represented by a PDA. Other constructions can also be represented by appropriate PDAs such as the caching schemes based on liner block code [21], projective space [26], and combinatorial designs [27], etc. We list their performances in Table I. In addition, given a PDA, we can obtain a PDA with any number of users by the following result.
@@ III Non-half-sum disjoint packing
Recall that Condition C3333-a) of Definition 1 is called the Latin property in combinatorial design theory. Naturally, we can construct a PDA with $K=F$ by first constructing a Latin square and then putting stars in some integer entries such that the resulting array satisfies conditions C1111, C2222, and C3333. We then obtain the desired PDA with linear subpacketization. Specifically, we use the classical and simple construction of the Latin square $\mathbf{L}=(l_{f,k})_{f,k\in\mathbb{Z}_{v}}$ where the entry $l_{f,k}=f+k$ for each $f$, $k\in\mathbb{Z}_{v}$. Recall that $\mathbb{Z}_{v}$ is a positive integer ring, i.e., the arithmetic operations in the ring $\mathbb{Z}_{v}$. When we cyclically replace the integer entries by stars, condition C3333 of Definition 1 is transformed into the following novel combinatorial structure for the integer entries of the first row of the Latin square.
@@
  • For each $\mathcal{D}\in\mathfrak{D}$, the half-sum of any two different elements in $\mathcal{D}$ (i.e., the sum of the two elements divided by 2222111The operations are in $\mathbb{Z}_{v}$ and since $v$ is odd, then $2$ has an inverse, that is, $1/2$ in $\mathbb{Z}_{v}$ is also an element of $\mathbb{Z}_{v}$) does not appear in any block of $\mathfrak{D}$ [^2] In an NHSDP $(\mathbb{Z}_{v},\mathfrak{D})$, each element of $\mathfrak{D}$ is called block.
@@
It is not difficult to check that each column of $\mathbf{P}$ has exactly $Z=v-bg=15-2\times 4=7$ stars, there are exactly $S=30$ vectors in $\mathbf{P}$, and each vector occurs at most once in each row and each column. So the conditions C1111, C2222, and C3333-a) of Definition 1 hold. Finally, let us consider condition C3333-b). Let us first consider the entries $p_{1,0}=p_{0,1}=(1,1)$. We have $p_{1,1}=p_{0,0}=*$ which satisfies condition C3333-b). The reason for this is as follows. When $f=1$, if $k=0$ we have $k-f=-1=14\in\mathcal{D}_{1}$ in (12) and $1+0=1$, then we set $p_{1,0}=(1,1)$; if $k=1$, we have $k-f=0$, which does not appear in either $\mathcal{D}_{1}$ or $\mathcal{D}_{2}$. Thus, we set $p_{1,1}=*$. Similarly, we can check that all the vectors in $\mathbf{P}$ satisfy condition C3333-b) of Definition 1. Therefore, $\mathbf{P}$ is a $(15,15,7,30)$ PDA that can realize a coded caching scheme with the memory ratio $\frac{M}{N}=\frac{7}{15}$, the subpacketization $F=K=15$, and load $R=2$. $\square$
In Example 2, each column of $\mathbf{P}$ has exactly $|\mathcal{D}_{1}|+|\mathcal{D}_{2}|=|\mathcal{D}_{1}\cup\mathcal{D}_{2}|=8$. This implies that condition C1111 of Definition 1 is guaranteed by the first condition of NHSDP. From (13), the half-sum set is =$\mathcal{H}=\{0,3,6,7,8,9,12\}$. We can see that in Example 2, the stars of $\mathbf{P}$ in (15) are exactly the entries $p_{i,i+h}$ for each $i\in\mathbb{Z}_{15}$ and $h\in\mathcal{H}$. So the half-sums of all the blocks generates the star positions. This implies that condition C3333-b) of Definition 1 is ensured by the second condition of NHSDP. Recall that our construction is based on the constructing a Latin square. So the condition C3333-a) always holds.
Then, for any parameters $v$, $g$ and $b$, if there exists a $(v,g,b)$ NHSDP by Construction 1, we can also obtain a PDA in the following theorem, whose proof is given in Appendix A.
@@ IV Constructions of PDAs via NHSDPs
For instance, the integer $31$ in $\mathcal{D}_{(1,1,1)}$ can be obtained by $1+5+25=31$. The number of all possible $(\alpha_{1},\alpha_{2},\alpha_{3})$ is $82^{3}=$ 8; thus the block $\mathcal{D}_{(1,1,1)}$ contains $8$ different integers. Similarly, we can obtain the following blocks.
$\begin{split}\mathfrak{D}=\{&\mathcal{D}_{(1,1,1)}=\{31,21,29,19,-19,-29,-21,-31\},\\ &\mathcal{D}_{(2,1,1)}=\{32,22,28,18,-18,-28,-22,-32\},\\ &\mathcal{D}_{(1,2,1)}=\{36,16,34,14,-14,-34,-16,-36\},\\ &\mathcal{D}_{(2,2,1)}=\{37,17,33,13,-13,-33,-17,-37\},\\ &\mathcal{D}_{(1,1,2)}=\{56,46,54,44,-44,-54,-46,-56\},\\ &\mathcal{D}_{(2,1,2)}=\{57,47,53,43,-43,-53,-47,-57\},\\ &\mathcal{D}_{(1,2,2)}=\{61,41,59,39,-39,-59,-41,-61\},\\ &\mathcal{D}_{(2,2,2)}=\{62,42,58,38,-38,-58,-42,-62\}\}.\end{split}$ (20)
@@
By the above equation, $\mathfrak{D}$ has $m_{1}m_{2}m_{3}=8$ blocks and the intersection of any blocks is empty. The minimum and maximum values of $\mathfrak{D}$ are $-62$ and $62$, respectively. To ensure that these two sets $[-62:0]$ and $[1:62]$ do not have overlap in $\mathbb{Z}_{v}$, the value of $v$ must satisfy $v\geq 2\times 62+1=125$, since 00 is also in $\mathbb{Z}_{v}$. Next, let we verify that $(\mathbb{Z}_{125},\mathfrak{D})$ is an $(125,8,8)$ NHSDP. Now let we consider the half-sum of $31$ and $21$, i.e., $\frac{31+21}{2}=26$, which is not in any block of $\mathfrak{D}$. Similarly we can check the half-sum of any two different integers in each block of $\mathfrak{D}$.
The intuition on why the proposed construction leads to a NHSDP is summarized as follows.
  • In order to ensure that the first condition of NHSDP, i.e., the intersection of any two blocks is empty, each integer in $\mathfrak{D}$ is uniquely represented by a linear combination of all elements in $\mathcal{X}$, and the coefficients of $(x_{1},x_{2},x_{3})$ are $(a_{1}\alpha_{1},a_{2}\alpha_{2},a_{3}\alpha_{3})\in\mathcal{C}=\{\pm 1,\pm 2\}^{3}$, since $(a_{1},a_{2},a_{3})\in[2]^{3}$ and $(\alpha_{1},\alpha_{2},\alpha_{3})\in\{-1,+1\}^{3}$. By the above structure of linear combinations and by the choice of $\{x_{i}:i\in[3]\}$, each block $\mathcal{D}_{\mathbf{a}}$ where $\mathbf{a}\in[2]^{3}$ has $82^{3}=$ 8 integers and any two different blocks $\mathcal{D}_{\mathbf{a}_{1}}$ and $\mathcal{D}_{\mathbf{a}_{2}}$ do not have any overlap.
  • We then focus on the second condition of NHSDP, i.e., the half-sum of any two different elements in one block does not appear in any block of $\mathfrak{D}$. For any two different integers in $\mathcal{D}_{{\bf a}}$, say
    $d=\alpha_{1}a_{1}x_{1}+\alpha_{2}a_{2}x_{2}+\alpha_{3}a_{3}x_{3}\ \ \text{and}\ \ d^{\prime}=\alpha^{\prime}_{1}a_{1}x_{1}+\alpha^{\prime}_{2}a_{2}x_{2}+\alpha^{\prime}_{3}a_{3}x_{3},$
    where $(\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3})\in\{-1,1\}^{3}$ and $(\alpha_{1},\alpha_{2},\alpha_{3})\neq(\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3})$, we have their half-sum
    $\frac{d+d^{\prime}}{2}=\frac{\alpha_{1}+\alpha^{\prime}_{1}}{2}a_{1}x_{1}+\frac{\alpha_{2}+\alpha^{\prime}_{2}}{2}a_{2}x_{2}+\frac{\alpha_{3}+\alpha^{\prime}_{3}}{2}a_{3}x_{3}.$
    Recall that $(\alpha_{1},\alpha_{2},\alpha_{3})\neq(\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3})$ there is at least one item of $\frac{\alpha_{1}+\alpha^{\prime}_{1}}{2}$, $\frac{\alpha_{2}+\alpha^{\prime}_{2}}{2}$ and $\frac{\alpha_{3}+\alpha^{\prime}_{3}}{2}$ equal to 00. This implies that each half-sum can only be represented by a linear combination of all elements in $\mathcal{X}$ with coefficients in $\mathcal{C}^{\prime}=[-2:2]^{3}\setminus\mathcal{C}$. Hence, $\mathcal{C}\cap\mathcal{C}^{\prime}=\emptyset$, which implies that each half-sum does not occur in $\mathcal{D}$.
@@
  • The WCWL scheme (see Table I): When $K=125$ and $t=61$ in [14] we have a WCWL scheme with $K=125$, the subpacketization $F=125$, and the transmission load $\frac{K-t}{2\lfloor\frac{K}{K-t+1}\rfloor}=\frac{64}{3}\approx 21.333>8$. Clearly our scheme has a lower transmission load while maintaining a lower memory ratio and the same subpacketization;
  • The CWZW scheme (see Table I): In the case of $q=2$ and $m=7$, the scheme in [12] includes the schemes in [9, 22, 21]. So we only need consider the CWZW scheme. First we can obtain a CWZW scheme with $K=16$, the memory ratio $M/N=0.5>\frac{61}{125}$, the subpacketization $F^{\prime}=2^{7}=128>125$, and the transmission load $1$. Using the grouping method in Lemma 2, we have a grouping CWZW scheme with $K=125$, $M/N=0.5>\frac{61}{125}$, the subpacketization $F_{1}=\frac{16}{\text{gcd}(125,16)}\cdot F^{\prime}=2048>125$, and the transmission load $\frac{125}{16}=7.8125$. We can see that our scheme has a lower memory ratio and a smaller subpacketization than the scheme in [12], albeit with a slight increase in the transmission load;
  • The MN scheme (see Table I): An exhaustive computer search shows that we can obtain a $(125,M,N)$ grouping MN coded caching scheme with $M/N=0.5$, an appropriate subpacketization (${10\choose 5}=252$, and the transmission load $R=\frac{5}{6}\times\frac{125}{10}=\frac{125}{12}>10>8$ by using a $(10,M,N)$ MN scheme based on the grouping method in Lemma 2. It is evident that all the memory ratio, the subpacketization, and the transmission load are larger than our proposed scheme.
@@
By Construction 2, for any $n$ and $m_{1}$, $m_{2}$, $\ldots$, $m_{n}$ are positive integers, suppose the vector $\mathbf{a}=(a_{1},a_{2},\ldots,a_{n})=(m_{1},m_{2},\ldots,m_{n})\in\mathcal{A}$. The corresponding block $\mathcal{D}_{\mathbf{a}}$ is given by
@@
which contains the integers $-f(1)-f(2)-\dots-f(n)$ and $f(1)+f(2)+\ldots+f(n)$. To ensure that $[-f(1)-f(2)-\dots-f(n):0]$ and $[1:f(1)+f(2)+\ldots+f(n)]$ can not overlap in $\mathbb{Z}_{v}$, the value of $v$ must satisfy
@@
Next, given the parameters $v$ and $n$ (recall that $K=v$ and $g=2^{n}$), we consider the selection of $m_{1},\ldots,m_{n}$, such that the memory ratio is minimum. In other words, fixing the coded caching gain, we aim to search the coded caching scheme requiring the minimum memory ratio. Considering the constraint $v\geq 2\phi(m_{1},m_{2},\ldots,m_{n})+1$ into the optimization problem for the selection of $m_{1},m_{2},\ldots,m_{n}$, we have
@@
Problem 1 is an integer programming problem, which is NP-hard. We first relax the constraint $m_{1},m_{2},\ldots,m_{n}\in\mathbb{Z}^{+}$ to $m_{1},m_{2},\ldots,m_{n}\in\mathbb{R}^{+}$, in order to simplify the problem to a convex optimization problem, which is then solved by the Lagrange Multiplier Method. Finally we take the floor operation to the solution. For the ease of notation, we define that $q:=K^{1/n}$. The following theorem introduces the sub-optimal solution by the above optimization strategy, whose proof is given in Appendix C.
@@
which can realize a coded caching scheme with the memory ratio $\frac{M}{N}=1-\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}}{q^{n}}$ and the transmission load $R={\left\lfloor\frac{q-1}{2}\right\rfloor}^{n}$.
@@
By Theorem 3, given a user number $K$ there are $\left\lfloor\log_{p}K\right\rfloor$ memory-rate ratio points for each positive $p\geq 3$. In addition, when $q$ is equal to a positive odd integer such that $q\geq 3$, the solution in (23) is an optimal solution to Problem 1. Then by Theorem 3 we have the following result.